OpenStudy (anonymous):

quick question : how would i do this? [(X+3)/2]^2 + [(y-4)/2]^2 = 1

OpenStudy (anonymous):

\[(\frac{ x+3 }{ 2 })^{2} +(\frac{ y-4 }{ 2 })^{2}=1\]

OpenStudy (anonymous):

If it's the centre of the circle you want, it's (-3/2,2). If it's the radius, it's 1. I don't get the question in this. :\

OpenStudy (anonymous):

i guess i just want to simplify it

OpenStudy (anonymous):

i dont know what to do with the exponent?

OpenStudy (anonymous):

Oh, eff the radius is 2.

OpenStudy (anonymous):

It's a circle, open up the damn thing!

OpenStudy (perl):

circles are technically ellipsi

OpenStudy (perl):

ellipses**

OpenStudy (anonymous):

okay, just help me simplify

OpenStudy (anonymous):

i know what needs to be done, just help me distribute that exponent

OpenStudy (anonymous):

\[x ^{2} + 6x+ 9 + y^{2} -8y + 16=4\]

OpenStudy (anonymous):

okay break it down for me

OpenStudy (anonymous):

what do i do first?

OpenStudy (anonymous):

how do i distribute that exponent?

OpenStudy (anonymous):

you know that property about [(a+b)^{2}] Right?

OpenStudy (anonymous):

but theres something right here it i woudl have to foil (a+b)(a+b)

OpenStudy (anonymous):

It's the same thing. It's happening in (x+3)^2 and (y-4)^2.

OpenStudy (anonymous):

a^2+2ab+b^2

OpenStudy (anonymous):

what about that 2 in the demominator?

OpenStudy (anonymous):

I'm taking thw two from both the denominators out.

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

Mmhmm y'can. See, what I did there was, (1/2)^{2}*(x+3)^2 + (1/2)^{2}*(y-4)^2=1 and then, after squaring (1/2) and taking it common in the LHS, I multiply both sides of the equation by 4.

OpenStudy (anonymous):

Which gave me (x+3)^2+(y-4)^2=4

OpenStudy (anonymous):

Hope that helped.

OpenStudy (anonymous):

i dont know if i would just multiply 1/2 to both sides... dont i have to stick to order of operations?

OpenStudy (anonymous):

exponents first...?

OpenStudy (anonymous):

P.E.M.D.A.S?

OpenStudy (anonymous):

You're making that complicated. Look, it's a valid term. You can simplify the exponents before solving the equation.

OpenStudy (anonymous):

are you sure? i seems like i would be bound to use the exponent first

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Even if we go by PEDMAS, the solution isn't different, just this approach is easier... and no problem, you can solve the exponent first. \[\left( \frac{ x+3 }{ 2 } \right)^{2} +\left( \frac{ y-4 }{ 2 } \right)^{2} = 1\] \[\left( \frac{ x^{2} }{ 4 } +\frac{ 6x }{ 4 }+ \frac{ 9 }{ 4 }\right) + \left( \frac{ y^{2} }{ 4 } -\frac{ 8y }{ 4 } +\frac{ 16 }{ 4 }\right)\\] = 1

OpenStudy (anonymous):

well i'll be... i will trust you and just multiply by 2

jhonyy9 (jhonyy9):

why by 2 and not by 4 ?

jhonyy9 (jhonyy9):

(x+3)^2 (y-4)^2 ------- + -------- = 1 2^2 2^2 take it on common denominator so for you need multiplie the right side by 4 than will get like result (x+3)^2 +(y-4)^2 =4 ok ?

jhonyy9 (jhonyy9):

can you continue it now ?

OpenStudy (anonymous):

I did it by multiplying by 4. -_-"